### v1

# Credit Card Fraud Detection Using Advanced Transformer Model

Abstract
With the proliferation of various online and mobile payment systems, credit card fraud has emerged as a significant threat to financial security. This study focuses on innovative applications of the latest Transformer models for more robust and precise fraud detection. To ensure the reliability of the data, we meticulously processed the data sources, balancing the dataset to address the issue of data sparsity significantly. We also selected highly correlated vectors to strengthen the training process.To guarantee the reliability and practicality of the new Transformer model, we conducted performance comparisons with several widely adopted models, including Support Vector Machine (SVM), Random Forest, Neural Network, and Logistic Regression. We rigorously compared these models using metrics such as Precision, Recall, and F1 Score. Through these detailed analyses and comparisons, we present to the readers a highly efficient and powerful anti-fraud mechanism with promising prospects. The results demonstrate that the Transformer model not only excels in traditional applications but also shows great potential in niche areas like fraud detection, offering a substantial advancement in the field.
Index Terms:
Credit card Fraud Detection, Transformer, preprocessing, Precision, Recall, F1-Score
I Introduction
The emergence of information technologies such as mobile internet and big data has led to the rise of internet finance, bringing convenience to people’s daily lives. However, along with this convenience comes the phenomenon of internet financial fraud, with credit card transactions being the primary target of various fraudulent activities. The rapid development of internet financial services has significantly increased the probability of credit card fraud incidents, resulting not only in immeasurable losses for individuals and businesses but also in numerous socio-economic problems.
A major technical challenge lies in the fact that fraudulent transactions account for only a small portion of credit card payments, making it challenging for machine learning algorithms to detect such sensitive behavior[16]. As the most powerful and capable model to date, the Transformer large model exhibits revolutionary identification capabilities, making it particularly suitable for recognizing fraudulent data in small samples.
The remainder of this paper is organized as follows. In Section II, we provide a comprehensive review of the related research that has informed the objectives and methodological choices of the current study. This review covers key contributions in areas such as statistical modeling, neural networks, and machine learning approaches to credit card fraud detection[21].
Section III then delves into the detailed methodology employed in this work, explaining the data processing, sampling strategies, model architectures, and underlying principles that constitute our fraud detection framework.
Moving to Section IV, we present the results of our extensive experimental evaluation, comparing the performance of our proposed model against several widely-used benchmark methods to rigorously assess the detection accuracy achieved.
Finally, Section V concludes the paper by summarizing the key findings, discussing the practical implications of our work, and outlining promising directions for future research in the field of credit card fraud detection.
II Related Work
The detection of fraud in financial transactions has long been a pressing research challenge in the field of payment systems. As the use of mobile payments, digital wallets, and web-based payment methods has proliferated, accurately identifying the small proportion of fraudulent transactions amidst the vast volume of payment data has become an increasingly critical issue.
The academic literature has accumulated a rich body of research addressing this problem. Bolton and Hand[1] pioneered the use of statistical models to analyze and detect the presence of fraudulent activity. Building upon this foundation, Stojanović[15] subsequently employed more advanced neural network models for fraud detection. Christoph[5] proposed a machine learning approach involving feature vectorization and classification, providing an experimental framework for related studies. Furthermore, the work of Weng et al[19]. offered valuable insights by demonstrating how data analysis and AI techniques can be leveraged to tackle data security and cybersecurity challenges, inspiring the focus of the current research.
In the realm of model selection, the success of large language models in other domains, as showcased by the research of Wang et al[17]., has provided important guidance. To ensure the efficacy of training and testing strategies, the data cleaning and processing methods developed by Wei et al[18]. have been invaluable. Shen et al[14]s work on efficient classification models has offered valuable experience in model selection, while Ding et al.’s[2] approach to data filtering, segmentation, and augmentation has provided useful insights for addressing sparse data challenges.
When it comes to experimental comparisons, the detailed evaluations of supervised learning, unsupervised learning, and deep learning methods conducted by Kazeem et al[4]. have served as a useful reference. Additionally, the rigorous experimental and analytical methodologies employed by Jha et al[3]. have informed the meticulous implementation of the current study.
This rich body of prior research has provided a solid foundation and has inspired novel ideas for our own exploration. Building upon these insights, we aim to contribute innovative methods that advance the state of the art in credit card fraud detection, offering more effective solutions for the industry.
III Methodology
In Section III, we provide a detailed account of the methodology employed in this work. This includes a thorough discussion of the data processing techniques utilized, such as ensuring an equally distributed dataset. We also examine the relevant prior research on feature engineering and dimensionality reduction, and analyze the impact these techniques had on the data used in our study. By comprehensively documenting these methodological components, we aim to establish transparency and enable the replication of our research approach.
The primary focus of our methodology lies in accurately identifying instances of credit card fraud in datasets where fraudulent samples are exceedingly rare. We conducted comprehensive evaluations and explorations of various model architectures, accompanied by detailed and thorough data analysis and preprocessing.[19] Additionally, we selected relevant loss functions and evaluation methods to ensure the robustness and effectiveness of our approach.
III-A Dataset Introduction
The dataset utilized in this study comprises European credit card transaction records, sampled from real data spanning two consecutive days in September 2013. Among the dataset’s 284,804 transactions, 492 were identified as fraudulent, accounting for 0.172% of the total. These data encompass multiple variables, denoted as v1 through v28, with their specific distributions depicted.
III-B Data Processing
We first resample and preprocess the dataset to address the issue of data imbalance. By randomly sampling and merging, a new balanced dataset is generated, containing an equal number of fraudulent and non-fraudulent transaction samples. The new data set is then shuffled to ensure the randomness and reliability of the samples.
III-B1 Equally Distributing and Correlating
To generate an equally distributed dataset, the initial step involves displaying textual output showcasing the proportions of class distribution to confirm the balance of the dataset[6]. Following this, a bar plot is generated utilizing Seaborn’s countplot function to visually depict the distribution of classes. This visualization illustrates an equal distribution of classes, thereby underscoring the effectiveness of the subsampling technique in achieving class balance, as shown in Figure 2. A careful comparison of Figure 1 and Figure 2 reveals that the issue of data sparsity has been effectively addressed[7].
Figure 1: Figure 1: Transaction Class Distribution
Figure 2: Figure 2: Equally Distributed
III-B2 Feature Correlation Analysi
In our analysis, we meticulously compared the original imbalanced dataset with its subsampled counterpart, unearthing a plethora of advantages stemming from data balancing[23]. This encompassed a noticeable boost in model performance, mitigated susceptibility to overfitting, heightened predictive stability, and augmented model interpretability. The correlation efficacy visibly escalated as evidenced by the comparison, underscoring the tangible benefits derived from this approach. Correspond info please check Figure 3 and Figure 4.
Figure 3: Figure 3: Imbalanced Correlation matrix
Figure 4: Figure 4: Subsample correlation matrix
III-B3 Dimensionality reduction
To achieve optimal dimensionality reduction for the data, three different techniques—T-SNE, PCA, and Truncated SVD—are employed to map the data onto a two-dimensional space[18]. This facilitates a visual representation of the distribution of data points on scatter plots, aiding in a better understanding of the data’s structure and clustering tendencies[22]. Compared to PCA and Truncated SVD, T-SNE excels in preserving the nonlinear structures and local relationships within high-dimensional data. Consequently, T-SNE offers superior advantages in data visualization and clustering analysis. The specific comparison results are illustrated in the following Figure 5.
Figure 5: Figure 5 T-SNE
Figure 6: Figure 6 PCA
Figure 7: Figure 7 Trucated SVD
III-C Model architectures
III-C1 Transformer Structure
The Transformer architecture employed in this study primarily consists of a Self-Attention Mechanism and a Feed-Forward Neural Network.
Self-Attention Mechanism
The Self-Attention Mechanism is utilized to compute attention weights for each position in the input sequence, capturing dependencies between different positions. In the Transformer Encoder[24], the Self-Attention Mechanism is referred to as Multi-Head Attention. Below is the mathematical formulation of Multi-Head Attention:
Given an input sequence $X=\{x_{1},x_{2},\ldots,x_{n}\}$, we first map the input sequence to Query ($Q$), Key ($K$), and Value ($V$) spaces. By matrix multiplication and scaling, we obtain Query ($Q$), Key ($K$), and Value ($V$) tensors:
$Q=XW_{Q}\quad K=XW_{K}\quad V=XW_{V}$
where $W_{Q}$, $W_{K}$, and $W_{V}$ are learnable weight matrices.
Next, we compute the attention weights ($A$) by taking the dot product of Query and Key, followed by scaling and softmax normalization:
$A=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)$
where $d_{k}$ is the dimensionality of Query and Key.
Finally, we multiply the attention weights with the values and then perform a weighted sum to obtain the output $Z$ of the Self-Attention Mechanism[9]:
$Z=AV$
The Multi-Head Attention mechanism calculates the outputs of multiple attention heads by using independent Query, Key, and Value projection matrices. These outputs are concatenated along the last dimension to obtain the final output of the Self-Attention Mechanism.
Operations in Transformer Encoder Layer
In the Transformer encoder layer, we apply the self-attention mechanism to the input sequence, followed by a fully connected feed-forward neural network.
In the self-attention mechanism, we use the input sequence as queries, keys, and values, and compute attention weights for each position with respect to other positions[11]. Then, these attention weights are used to weight-sum the values of the input sequence, resulting in the output of the self-attention mechanism[20].
Next, the output of the self-attention mechanism is transformed through a feed-forward neural network, and the result is added back to the input (using residual connections)[12], followed by layer normalization. Mathematically, it can be represented as:
$\text{EncoderLayer}(X)$ $=\text{LayerNorm}(X+\text{Attention}(X,X,X))$
$+\text{FeedForward}(X)$
Here, $X$ represents the input sequence, Attention represents the self-attention mechanism, and FeedForward represents the feed-forward neural network.
Here is the structure of our implemented model Figure 6:
Figure 8: Figure 8: Transformer Structure[13]
IV Evaluation
In this section, we first introduce the evaluation metrics, and then present our experiment results.
IV-A Evaluation Metrics
In this experiment, our evaluation metrics include Precision, Recall, and F1-score. These metrics provide different perspectives when assessing the classification results of the model[10].
Precision
Precision refers to the proportion of true positive samples among all samples predicted as positive by the model. It can be represented by the following formula:
$\text{Precision}=\frac{TP}{TP+FP}$
where TP (True Positives) is the number of true positive samples, and FP (False Positives) is the number of false positive samples.
Recall
Recall indicates the proportion of true positive samples among all true positive samples. It can be represented by the following formula:
$\text{Recall}=\frac{TP}{TP+FN}$
where TP (True Positives) is the number of true positive samples, and FN (False Negatives) is the number of false negative samples.
F1-score
F1-score is the harmonic mean of precision and recall, providing a balanced measure that combines both metrics’ performance. It can be represented by the following formula:
[…]
The choice of these metrics depends on the specific problem and requirements. For example, in some cases, there may be a greater emphasis on avoiding false negatives, in which case recall might be prioritized. In other cases, precision might be more important. The F1-score provides a comprehensive evaluation that considers both precision and recall.
IV-B Results
In pursuit of fairness and precision in selecting the most outstanding fraud detection model, we meticulously curated a variety of different classification models[8]. Among the designs were Logistic Regression, K-Nearest Neighbor, Support Vector Machine, Decision Tree Classifier, and Neural Network[14]. Employing an identical dataset, we subjected these models to rigorous comparison to discern their efficacy.
To obtain the most precise data, we have implemented data balancing techniques and then randomly sampled both fraudulent and non-fraudulent samples. We have conducted a detailed comparative analysis, including evaluations of F1-score, recall, and precision, to assess the accuracy of the models. The results have been consolidated, considering multiple case scenarios, to arrive at a comprehensive set of parameters.
TABLE I: Test Results
Model Precision Recall F1-score
KNN 0.93 0.92 0.92
SVM 0.93 0.93 0.93
Decision Tree 0.91 0.90 0.89
Logistic Regression 0.96 0.96 0.96
Neural Network 0.975 0.999 0.988
Transformer 0.998 0.998 0.998
Interestingly, the Transformer-based model has demonstrated outstanding performance across the various data comparisons, exhibiting no inferiority even when pitted against the high-performing Neural Network model. Given the rapid advancements of Transformer models in the field of Generative AI, as exemplified by the success of ChatGPT, the Transformer model also exhibits immense potential in the specialized domain of fraud detection.
V Conclusion
In this experiment, our greatest challenge lay in adapting the Transformer, originally designed for natural language processing tasks, into a classification model suitable for fraud detection. Over several months, we devoted considerable effort to understanding the underlying principles and fine-tuning the Transformer’s Encoder layers for classification purposes. The resulting model lived up to expectations, demonstrating exceptional accuracy across F1-score, precision, and recall metrics. The outcome of this experiment serves to underscore two key points. Firstly, it validates that Transformer models excel not only in general artificial intelligence applications like ChatGPT but also possess significant potential in the realm of classifier models when compared to traditional, classical approaches[17]. Secondly, our innovative application of the Transformer in fraud detection lays a solid foundation for future development of optimized Transformer-based models, promising enhanced financial security measures.